# Reciprocal relation of Schwinger pair production between $\textrm{dS}_{2}$ and $\textrm{AdS}_{2}$

### v1

@@ III.1 Charged Scalar Production in AdS2
Imposing the boundary condition $\phi_{\mathrm{out}}^{(-)}=0$ that describes the tunneling process
@@
the Bogoliubov coefficients $\alpha$ and $\beta$ are defined as [24]
@@
and the vacuum persistence amplitude square and the Bogliubov relation
@@ III.2 Charged Spinor Production in AdS2
The BF bound for the stability of AdS space is given by $T_{\rm U}\leq T_{\rm GH}$, i.e., $|qE|\leq mH$. A physical meaning is that the circular motion of charge $q$ by the electric field in the Euclidean time extends over the AdS space. Pair production requires circular loops to reside inside the AdS radius, and thus the Schwinger pair production violates the BF bound.
@@ IV Charged Spinor Production in Planar (A)dS Coordinates
The planar coordinates for the ${\rm AdS}_{2}$ space
@@
When $H\rightarrow 0$, the metric tensors and potentials have the limit of the Minkowski space, i.e. $A=-Etdx$ and $A=Exdt$, respectively.
@@ VI Phase-Integral Formulation of Pair Production
[…]
@@
[…]
@@
where the first term comes from a contour excluding both ±$\pm i$ and the second and third terms come from contours enclosing only one pole $i$ or $-i$, and the last term from the contour enclosing both poles ±$\pm i$. The result is consistent with the common leading term for scalars (24) and spinor (41).
@@
[…]
@@ VII Conclusion
We have exactly obtained the Schwinger production of spinors (spin-1/2 fermions) and scalars (spin-0 bosons) in a uniform electric field in the global coordinates of (A)dS2 and compared the results with those in the planar coordinates. The Dirac and Klein-Gordon equations in a uniform electric field respect the symmetry SO(2,1) (SO(1,2)) in (A)dS2 and have solutions in terms of hypergeometric functions, from which the Bogoliubov relations are found between the adiabatic vacua in two asymptotic regions and the mean numbers of Schwinger pair production of spinors and scalars are calculated in the in-out formalism. In the global dS2 space, for instance, the spinor production rate per unit two volume has an invariant expression
@@
where ${\cal F}=E^{2}/2$ is the Maxwell scalar, and $R=2H^{2}$ the scalar curvature. Here ${\cal D}=g(q\sqrt{2{\cal F}}/2\pi)\times\sqrt{1+m^{2}R/4q^{2}{\cal F}}$ is the density of states with the spin multiplicity ($g=2$). The formulas recover the limits both of Minkowski space with the electric field and pure dS2 space without any electric field.
First, in the Minkowski limit ($R=2H^{2}=0$), the spinor production rate recovers the Schwinger formula in two dimensions:
@@
Second, in the limit of zero electric field or charge neutral particle, the spinor production rate recovers the pure dS2 in global coordinates [35] except for the density of states ${\cal D}=gmH/2\pi$, which is needed to find the QED effective action for pure dS2. However, there is no production of spinors in the AdS2 without electric field, and furthermore, the BF bound $qE/H^{2}\leq m/H$ prohibits weak electric fields from producing any spinor.
The main results show the combined effect of an electric field and spacetime curvature for pair production in a nontrivial way. The crucial factor is the ratio of the Maxwell scalar to the scalar curvature square or the ratio of the dS radius to the electric length ($l_{\rm E}=1/\sqrt{qE}$), as shown in (108). Another factor is the ratio of the dS radius to the Compton wavelength of charge. Except for the two limits of the electric field or the curvature dominance, the QED and curvature effects are strongly coupled and intertwined in producing pairs. An intuitive way to understand the emission formula is to use the relevant temperatures: the Unruh temperature for charge acceleration by the electric field and the Gibbons-Hawking temperature in dS space. Then, the mean number has the Boltzmann factor ${\cal N}^{\rm(sp)}_{\rm dS}=\mathrm{e}^{-(m/T_{\rm GH})\times n_{\rm S}}$, in which the Gibbons-Hawking radiation is enhanced by an exponential folding factor $n_{\rm S}=1/(\sqrt{1+(T_{\rm U}/T_{\rm GH})^{2}}+T_{\rm U}/T_{\rm GH})$. When $T_{\rm GH}\gg T_{\rm U}$, the Boltmann factor may be interpreted as Gibbons-Hawking radiation of charge with a chemical potential $qEl_{\rm dS}$. In the opposite case of $T_{\rm U}\gg T_{\rm GH}$ when the Schwinger effect dominates over the Gibbons-Hawking radiation, the Boltzmann factor may be written as ${\cal N}^{\rm(sp)}_{\rm dS}=\mathrm{e}^{-(m/2T_{\rm U})\times n_{\rm GH}}$, in which the e-folding factor $n_{\rm GH}=2/(\sqrt{1+(T_{\rm GH}/T_{\rm U})^{2}}+1)$ gives the curvature effect on Schwinger pair production. This is a novel feature of strong coupling of gauge field and spacetime curvature.
@@ Acknowledgements.
The work of C.M.C. was supported by the National Science Council of the R.O.C. (Taiwan) under the grant NSTC 113-2112-M-008-027.

### v2

@@ III.1 Charged Scalar Production in AdS2
Imposing the boundary condition $\phi_{\mathrm{out}}^{(-)}=0$ that describes the tunneling process [^4] As will be shown in Sec. VI, the canonical space-dependent form (105) for spinless scalar describes a tunneling process with energy $(m^{2}/H^{2}+1/4)$ through a positive peaked potential barrier.
@@
the Bogoliubov coefficients $\alpha$ and $\beta$ are defined as [24], with normalized $\phi_{\mathrm{in}}^{(+)},\phi_{\mathrm{in}}^{(-)}$ and $\phi_{\mathrm{out}}^{(+)}$
@@
and the vacuum persistence amplitude square and the Bogoliubov relation
@@ III.2 Charged Spinor Production in AdS2
The BF bound for the stability of AdS space is given by $T_{\rm U}\leq T_{\rm GH}$, i.e., $|qE|\leq mH$. A physical meaning is that the circular motion of charge $q$ by the electric field in the Euclidean time extends over the AdS space. Pair production requires circular loops to reside inside the AdS radius, and thus the Schwinger pair production violates the BF bound. The e-folding factor for the Unruh effect is greater than one and suppresses the Schwinger effect.
@@ IV Charged Spinor Production in Planar (A)dS Coordinates
In order to compare the Schwinger effect in the global coordinates with that in the planar coordinates, we revisit Schwinger pair production of spinors in the planar coordinates of ${\rm dS}_{2}$ and, for the first time, study Schwinger pair production of spinors in the planar coordinates of ${\rm AdS}_{2}$. Then, we will show the relation of the Schwinger pair production between the planar ${\rm dS}_{2}$ and ${\rm AdS}_{2}$.
@@
Under time reversal $t\rightarrow-t$, the metric (76) maps to the other patch of the global Penrose diagram [5]. The planar coordinates for the ${\rm AdS}_{2}$ space
@@
When $H\rightarrow 0$, the metric tensors and potentials have the limit of the Minkowski space, i.e. A=EtdxA=-Etdx and $A=Exdt$, respectively.
@@ VI Phase-Integral Formulation of Pair Production
in which $\bar{\kappa}=\kappa+i\sigma/2$ and $\sigma=1,-1$ for spinor $\uparrow,\downarrow$, $\sigma=0$ for scalar, and $k$ is a half-integer for spinor and an integer for scalar. The phase of the positive frequency solution of (98) is given by $\tilde{\phi}_{k\sigma}(\tau)=\exp\left(-i\int d\tau\omega_{k\sigma}(\tau)\right)$. In the phase-integral formulation the leading mean number [34] is given by
@@
$\oint\frac{dz}{z^{2}+1}\sqrt{\mu^{2}(z^{2}+1)+2k\bar{\kappa}z-\Bigl(\bar{\kappa}^{2}-k^{2}+1/4\Bigr)}=\oint dz\frac{\mu}{z^{2}+1}\sqrt{(z-z_{1})(z-z_{2})},$ (100)
@@
where the first term comes from a contour excluding both $\pm i$ and the second and third terms come from contours enclosing only one pole $i$ or i-i, and the last term from the contour enclosing both poles $\pm i$. The result is consistent with the common leading term for scalars (24) and spinor (41).
@@
The phase $\tilde{\phi}_{\omega\sigma}(\zeta)=\exp\left(i\int d\zeta k_{\omega\sigma}(\zeta)\right)$ decays due to residues from simple poles at $z=\sinh(\zeta)=\pm i$ and another pole $z=\infty$ in the global coordinates and simple poles at $z=\exp(-\zeta)=0$ and $z=\infty$ in the planar coordinates. These poles contribute to the Boltzmann factor ${\cal N}_{\rm AdS}\simeq\mathrm{e}^{-2\pi(\kappa-\mu)}$ of the mean number of spinors or scalars in the global coordinates and ${\cal N}_{\rm AdS}\simeq\mathrm{e}^{2\pi(\mu-\mathrm{sgn}(\bar{\omega})\kappa)}$ in the planar coordinates.
@@ VII Conclusion
We have exactly obtained the Schwinger production of spinors (spin-1/2 fermions) and scalars (spin-0 bosons) in a uniform electric field in the global coordinates of (A)dS2 and compared the results with those in the planar coordinates. The Dirac and Klein-Gordon equations in a uniform electric field respect the symmetry SO(2,1) (SO(1,2)) in (A)dS2 and have solutions in terms of hypergeometric functions, from which the Bogoliubov relations are found between the in- and out-vacua with asymptotic WKB (adiabatic) in- and out-states, and the mean numbers of Schwinger pair production of spinors and scalars are calculated in the in-out formalism. In the global dS2 space, for instance, the mean number (41) gives rise to the spinor production rate per unit two volume, which may have an invariant expression in terms of the Maxwell scalar, ${\cal F}=E^{2}/2$, and the scalar curvature, $R=2H^{2}$,
@@
where ${\cal D}=g(H^{2}/2\pi)\mu=g(q\sqrt{2{\cal F}}/2\pi)\times\sqrt{1+m^{2}R/4q^{2}{\cal F}}$ is the density of states with the spin multiplicity ($g=2$) in an electric field that is uniformly distributed in the global geometry [56]. The formula recovers the limits both of Minkowski space with the electric field and pure dS2 space without any electric field.
First, in the Minkowski limit ($R=2H^{2}=0$), the spinor production rate recovers the Schwinger formula in two dimensions [57]
@@
Second, in the limit of zero electric field or charge neutral particle, the spinor production rate recovers the pure dS2 in global coordinates [35] except for the density of states ${\cal D}=gmH/2\pi$, which is required for the effective action in pure dS2. However, there is no production of spinors in the AdS2 without electric field, and furthermore, the BF bound $qE/H^{2}\leq m/H$ prohibits weak electric fields from producing any spinor.
The main results of this paper show the combined effect of an electric field and spacetime curvature for pair production in a nontrivial way. The crucial factor is the ratio of the Maxwell scalar to the scalar curvature square or the ratio of the dS radius to the electric length ($l_{\rm E}=1/\sqrt{qE}$), as shown in (108). Another factor is the ratio of the dS radius to the local Compton wavelength of charge ($1/m$). Except for the two limits of the electric field or the curvature dominance, the QED and curvature effects are strongly coupled and intertwined in producing pairs. An intuitive way to understand the emission formula is to use the relevant temperatures: the Unruh temperature for charge acceleration by the electric field and the Gibbons-Hawking temperature in dS space. Provided that the condition $l_{\rm E}<l_{\rm dS}$ for pair production is satisfied, the mean number (44) has the leading Boltzmann factor
${\cal N}^{\rm(sp)}_{\rm dS}=\mathrm{e}^{-(m/T_{\rm GH})\times n_{\rm S}},\quad n_{\rm S}=\frac{1}{\sqrt{1+(T_{\rm U}/T_{\rm GH})^{2}}+T_{\rm U}/T_{\rm GH}},$ (110)
in which the Gibbons-Hawking radiation is enhanced by an exponential folding factor:
$n_{\rm S}=\begin{cases}1-\frac{T_{\rm U}}{T_{\rm GH}}+\frac{1}{2}\Bigl(\frac{T_{\rm U}}{T_{\rm GH}}\Bigr)^{2}-\frac{1}{8}\Bigl(\frac{T_{\rm U}}{T_{\rm GH}}\Bigr)^{4}+{\cal O}\Bigl(\frac{T_{\rm U}}{T_{\rm GH}}\Bigr)^{6},\quad&(T_{\rm GH}>T_{\rm U}),\\ \frac{1}{2}\frac{T_{\rm GH}}{T_{\rm U}}-\frac{1}{8}\Bigl(\frac{T_{\rm GH}}{T_{\rm U}}\Bigr)^{3}+\frac{1}{16}\Bigl(\frac{T_{\rm GH}}{T_{\rm U}}\Bigr)^{5}+{\cal O}\Bigl(\frac{T_{\rm U}}{T_{\rm GH}}\Bigr)^{7},\quad&(T_{\rm U}>T_{\rm GH}).\end{cases}$ (111)
When $T_{\rm GH}\gg T_{\rm U}$, the Boltmann factor may be further interpreted as Gibbons-Hawking radiation of massive charge with a chemical potential $qEl_{\rm dS}$. In the opposite case of $T_{\rm U}\gg T_{\rm GH}$ when the Schwinger effect dominates over the Gibbons-Hawking radiation, the Boltzmann factor may be written as ${\cal N}^{\rm(sp)}_{\rm dS}=\mathrm{e}^{-(m/2T_{\rm U})\times n_{\rm GH}}$, in which the e-folding factor $n_{\rm GH}=2/(\sqrt{1+(T_{\rm GH}/T_{\rm U})^{2}}+1)$ gives the curvature effect on Schwinger pair production. This is a novel feature of strong coupling of gauge field and spacetime curvature, which may affect the Schwinger pair production in the early universe.
@@ Acknowledgements.
C. M. C. and S. P. K. thank the participants of the joint program [APCTP-2025-J01] held at APCTP, Pohang, Korea for fruitful discussions. The work of C. M. C. was supported by the National Science Council of the R.O.C. (Taiwan) under the grant NSTC 113-2112-M-008-027. The work of S. P. K. was supported in part by the Institute of Basic Science (Grant No. IBSR038-D1).